# Out-of-Distribution Detection: A Task-Oriented Survey of Recent Advances

### v3

@@ 6.3.2. Natural Language Processing
  • Intent Detection. Intent Detection is a significant application of OOD detection in NLP, particularly in dialogue systems, where identifying user intentions that fall outside the predefined set of intents is crucial (Wang et al., 2024c; Zhan et al., 2022; Akbari et al., 2025). Common datasets for evaluation include CLINC150 (Larson et al., 2019), Banking77 (Casanueva et al., 2020), and HWU64 (Liu et al., 2020b). These datasets are specifically designed to evaluate OOD detection in intent classification tasks.
  • Text Classification. In text classification tasks, OOD detection helps identify texts that do not belong to any of the known categories. Datasets such as News Category (Misra, 2022) and SST-2 (Socher et al., 2013) are frequently used for in-distribution data, with 20 Newsgroups (20NG) (Lang, 1995) often serving as an OOD dataset. Methods like VI-OOD (Zhan et al., 2024) enhance detection by optimizing joint distributions.
  • Question-Answering Systems. Question-answering systems require robust mechanisms to handle diverse user queries while maintaining response accuracy. The challenge of identifying questions that fall outside the system’s knowledge domain has led to sophisticated OOD detection approaches. Using standard in-distribution datasets (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), PubMedQA (Jin et al., 2019)) and OOD datasets (SQUAD 2.0 (Rajpurkar et al., 2018), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2019)), recent LLM-based methods (Zhang et al., 2024c) leverage likelihood ratios between pretrained and fine-tuned models for effective OOD detection.
  • Mathematical Reasoning. Mathematical problem-solving systems face challenges when encountering problem types not seen during training. Using MultiArith (Roy et al., 2016) as the primary in-distribution dataset, researchers evaluate OOD detection against diverse benchmarks including MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), SVAMP (Patil et al., 2021), AddSub (Hill et al., 2015), SingleEq (Wang et al., 2024b), and SingleOp (Wang et al., 2024b). Recent advances, such as TV-Score (Wang et al., 2024b), analyze embedding trajectory variations to enhance OOD detection performance.
@@ 6.3.3. Beyond Computer Vision and Natural Language Processing
  • Audio Recognition. In the field of audio recognition, OOD detection is essential for identifying inputs such as synthesized speech and dialects that diverge from the training distribution. Traditional probabilistic scoring methods often underperform on atypical audio samples (Lane et al., 2006), motivating the adoption of alternative approaches including confidence-based topic classification (Lane et al., 2006), neural sentence embeddings (Ryu et al., 2017), and deep nearest neighbor analysis (Bukhsh and Saeed, 2023). Autoencoder architectures built upon WavLM (Du et al., 2024b) have demonstrated improved performance, particularly in detecting synthetic speech. Standard datasets for evaluating OOD detection in this domain include MSCW (Mazumder et al., 2021) and Vocalsound (Gong et al., 2022), which provide both in-distribution and out-of-distribution samples for comprehensive benchmarking.
@@
  • Graph Neural Networks. Within graph neural networks (GNNs), OOD detection encompasses both node-level and graph-level tasks. Commonly explored techniques include uncertainty-based graph classification (Bazhenov et al., 2022), energy-based detection (Wu et al., 2023b; Jiang et al., 2025), and semi-supervised learning (Huang et al., 2022b). The HGOE framework (Junwei et al., 2024) enhances graph-level detection by incorporating anomalous samples during training, while PGR-MOOD (Shen et al., 2024) utilizes diffusion models to identify out-of-distribution molecular structures. GOLD (Wang et al., 2025b) generates pseudo-OOD data with significant differences from the ID embeddings by using a latent generative model to mimic the embeddings of training data and training it jointly with a GNN encoder and an OOD detector.Benchmarking typically relies on paired datasets from TU (Morris et al., 2020) and OGB (Hu et al., 2020), which facilitate systematic evaluation of OOD detection methods in GNNs.
  • Reinforcement Learning. OOD detection plays a crucial role in reinforcement learning (RL) by enhancing system robustness when encountering unprecedented states or actions (Leonardos et al., 2023). Common approaches include uncertainty quantification through dropout techniques and bootstrapping methods (Sedlmeier et al., 2019), as well as systematic environment modifications (Mohammed and Valdenegro-Toro, 2021). In the context of offline RL, advanced methods such as ADAC (Chen et al., 2025b) systematically assess OOD actions to effectively address distribution shifts. Comprehensive evaluation frameworks for OOD detection in RL primarily concentrate on various forms of environmental perturbations (Mohammed and Valdenegro-Toro, 2021), providing standardized testing scenarios for robustness assessment.
  • Time Series Analysis. In the domain of time series analysis, the inherent temporal dependencies present significant challenges for conventional OOD detection approaches. SRS (Belkhouja et al., 2023) addresses this by leveraging time series decomposition in conjunction with deep generative models to estimate conditional likelihoods, while CODiT (Kaur et al., 2023) introduces temporal equivariance to better capture sequential patterns. Recent studies have further evaluated modality-agnostic algorithms on multivariate time series datasets, outlining persistent challenges and future research directions (Gungor et al., 2025). In practice, OOD detection has been widely adopted in fraud detection services, such as the AWS Fraud Detector (web, ), which utilizes machine learning to identify transaction patterns that deviate from historical norms, effectively flagging potentially fraudulent activities in real time.
@@
  • Biological Classification. In biological classification, OOD detection plays a vital role in identifying previously unseen species or anomalies that extend beyond the categories present in training data. The integration of DNA barcodes has been shown to enhance image-based classification models through re-ranking strategies, thereby improving robustness and detection accuracy (Arslan et al., 2024).
@@
  • Spiking Neural Networks. Spiking Neural Networks (SNNs), recognized for their event-driven efficiency, present distinctive challenges for OOD detection due to their sparse latent representations. Recent methods have proposed combining feedforward learning with distance-based scoring and gradient-free attribution to address these challenges (Terres-Escudero et al., ). The N-MNIST dataset (Orchard et al., 2015) is commonly used for evaluation, with OOD samples generated through diverse event patterns.
  • Astronomical Imaging. In astronomical imaging, OOD detection is essential for distinguishing between simulated and real galaxy images, where conventional statistical techniques often prove inadequate. Approaches such as Bayesian model comparison and sparse variational autoencoders (Zhou et al., 2024) have been utilized to uncover subtle discrepancies. The SDSS dataset is widely adopted for benchmarking, with simulation-induced biases serving as OOD samples.
@@
  • Edge Computing. Within edge computing, efficient and privacy-preserving OOD detection is critical for devices with limited computational resources. SecDOOD (Li et al., 2025) addresses these constraints by employing a collaborative cloud-device architecture and HyperNetwork-based parameter generation, thereby eliminating the need for device-side backpropagation. CIFAR-10 and Tiny ImageNet are standard datasets for in-distribution and OOD evaluation in this context.
@@ 7.1. Better methodologies of OOD detection
Meta-Learning Adaptation. Rapid adaptation to novel data distributions at test time remains a significant challenge in OOD detection. Meta-learning algorithms, which follow a learning-to-learn paradigm, have emerged as promising solutions for enabling efficient adaptation to unseen test data. Additionally, the vast sample space of potential outliers in training-driven OOD detection motivates the development of improved sampling strategies for more effective utilization of outlier information (Ming et al., 2022a).
Model Selection. Selecting the most suitable OOD detection approach for a given task is non-trivial, as detection performance often varies across data distributions. Automated model selection frameworks, such as MetaOOD (Qin et al., 2025), leverage meta-learning to facilitate zero-shot model selection, while DSDE (Geng et al., 2024) and score-combining methods (Reehorst and Schniter, 2025) utilize proportion estimation and likelihood ratio tests to enhance ensemble and score fusion strategies. Interestingly, recent work has shown that partially trained, immature models can exhibit effective OOD detection capabilities (Montazeran and Köthe, 2025), suggesting new directions for leveraging model dynamics in detection.
Active Learning. Active learning strategies, which prioritize labeling the most informative data points, offer efficiency gains in data-scarce settings. In OOD detection, active learning can focus annotation efforts on ambiguous or high-uncertainty samples, thereby improving the model’s discrimination between in-distribution and out-of-distribution data. SISOM (Schmidt et al., 2024) exemplifies a unified approach that combines active learning and OOD detection by selecting samples based on feature space distances.
Continual and Incremental Learning. As real-world data is often encountered in a streaming and evolving manner, continual and incremental learning paradigms are increasingly relevant for OOD detection. These approaches enable models to adapt to new tasks and data distributions over time while retaining knowledge of previous ones. Recent efforts include continual unsupervised OOD detection (Doorenbos et al., 2024), systematic evaluation benchmarks like OpenCIL (Miao et al., 2024b), and hierarchical two-sample tests for continual OOD detection (Liu et al., 2025d).
@@ 7.2. More practical scenarios of OOD detection
Federated Environments. In practical scenarios, data is often distributed across multiple clients with heterogeneous distributions. Federated learning (FL) enables collaborative model training while preserving privacy, but also brings unique OOD detection challenges due to covariate and semantic shifts among clients. Addressing these challenges is crucial for robust federated models. For instance, FOOGD (Liao et al., 2024) combines score matching and feature invariance to handle simultaneous OOD shifts in federated settings.
Multi-Modal Detection. The proliferation of multi-modal data—ranging from images and text to audio and video—demands OOD detection methods capable of handling diverse inputs. Benchmarks such as MultiOOD (Dong et al., 2024) provide comprehensive testbeds for multi-modal OOD detection, introducing algorithms like A2D and NP-Mix for challenging scenarios. Further advancements, including DIAEF (Gao et al., 2024) and DPU (Li et al., 2024d), focus on cross-modal alignment and dynamic prototype updating, respectively. The integration of large language models (LLMs) into multi-modal OOD detection is poised to further enhance the field by combining advanced linguistic reasoning with other modalities.
Multi-Label OOD Detection. Real-world applications often involve multi-label data, where each instance may belong to multiple categories. Detecting OOD samples in these scenarios is challenging due to increased dimensionality and label correlations. Recent approaches (Mei et al., 2024; Aguilar et al., 2025) leverage uncertainty modeling and subjective logic to improve robustness in multi-label OOD detection.
Open-Vocabulary Scenario. Traditional approaches often assume access to all in-distribution categories, an assumption that does not hold in open-vocabulary contexts. Recent work (Li et al., 2024c) explores learning migratable negative prompts, enabling OOD detection even when only a subset of in-distribution labels is available, thus paving the way for more generalizable models.
@@
Long-Sequence Processing. In NLP, the detection of OOD inputs becomes increasingly complex when language models process sequences that exceed their pre-trained context window. In such scenarios, models frequently encounter unfamiliar dependency structures, increased attention demands, and altered positional encoding behaviors, all of which deviate from patterns seen during training. These distributional shifts impede the model’s generalization capabilities, highlighting the importance of thoroughly understanding their effects. Advancing OOD detection in this context is essential for ensuring reliable model performance on sequences that extend beyond standard context boundaries (Liu et al., 2025c).
@@ 7.3. New applications of OOD detection
Human-in-the-Loop Systems. The integration of human expertise into OOD detection is emerging as a critical direction, especially in high-stakes or safety-critical applications. Recent frameworks (Vishwakarma et al., 2024; Bai et al., 2024b, c) incorporate human feedback to refine detection thresholds and guide annotation, balancing algorithmic precision with human intuition. The AHA framework, for instance, strategically targets ambiguous regions for selective annotation, improving generalization and error correction under limited supervision.
Web Image Scraping. OOD detection is also finding novel applications in automating web image scraping. For example, zero-shot ID detection (Miyai et al., 2023c) classifies images as in-distribution if they contain relevant objects, offering a fresh perspective on leveraging OOD detection for large-scale, real-world data acquisition.
@@ 9. Acknowledgements
We sincerely appreciate the constructive comments and suggestions provided by all the reviewers and the editor, which have greatly improved the quality of our work. We are truly grateful for their valuable time and insightful feedback.

### v4

@@ 6.3.2. Natural Language Processing
  • Intent Detection. Intent Detection is a significant application of OOD detection in NLP, particularly in dialogue systems, where identifying user intentions that fall outside the predefined set of intents is crucial (Wang et al., 2024c; Zhan et al., 2022). Common datasets for evaluation include CLINC150 (Larson et al., 2019) and HWU64 (Liu et al., 2020b). These datasets are specifically designed to evaluate OOD detection in intent classification tasks.
  • Text Classification. In text classification tasks, OOD detection helps identify texts that do not belong to any of the known categories. Datasets such as News Category (Misra, 2022) and SST-2 (Socher et al., 2013) are frequently used for ID data, with 20 Newsgroups (20NG) (Lang, 1995) often serving as an OOD dataset. Methods like VI-OOD (Zhan et al., 2024) enhance detection by optimizing joint distributions.
  • Question-Answering Systems. Question-answering systems require robust mechanisms to handle diverse user queries while maintaining response accuracy. The challenge of identifying questions that fall outside the system’s knowledge domain has led to sophisticated OOD detection approaches.Recent LLM-based methods (Zhang et al., 2024c) leverage likelihood ratios between pretrained and fine-tuned models for effective OOD detection.
  • Mathematical Reasoning. Mathematical problem-solving systems face challenges when encountering problem types not seen during training. Using MultiArith (Roy et al., 2016) as the primary ID dataset, researchers evaluate OOD detection against diverse benchmarks. Recent advances, such as TV-Score (Wang et al., 2024b), analyze embedding trajectory variations to enhance OOD detection performance.
@@ 6.3.3. Beyond Computer Vision and Natural Language Processing
  • Audio Recognition. OOD detection in audio recognition has evolved from traditional probabilistic scoring (Lane et al., 2006) to methods leveraging confidence-based classification (Lane et al., 2006), neural embeddings (Ryu et al., 2017), and deep nearest neighbor approaches (Bukhsh and Saeed, 2023). Recent advances include autoencoder models based on WavLM (Du et al., 2024b), which are particularly effective for synthetic speech detection. Standard benchmarks such as MSCW (Mazumder et al., 2021) support comprehensive evaluation in this domain.
@@
  • Graph Neural Networks. In GNNs, OOD detection encompasses node- and graph-level tasks, utilizing uncertainty estimation (Bazhenov et al., 2022), energy-based methods (Wu et al., 2023b), and semi-supervised learning (Huang et al., 2022b). Frameworks like HGOE (Junwei et al., 2024) and PGR-MOOD (Shen et al., 2024) introduce anomalous or pseudo-OOD samples to enhance detection. GOLD (Wang et al., 2025b) further advances this area by generating pseudo-OOD data through latent generative modeling. Evaluation commonly relies on datasets from TU (Morris et al., 2020) and OGB (Hu et al., 2020).
  • Reinforcement Learning. OOD detection is vital in reinforcement learning for improving robustness to novel states and actions (Leonardos et al., 2023). Approaches typically involve uncertainty quantification (Sedlmeier et al., 2019), environment modifications (Mohammed and Valdenegro-Toro, 2021), and offline RL techniques such as ADAC (Chen et al., 2025b). Evaluation frameworks often focus on environmental perturbations (Mohammed and Valdenegro-Toro, 2021) to benchmark robustness.
  • Time Series Analysis. Time series OOD detection faces unique challenges due to temporal dependencies. Recent research explores modality-agnostic approaches for multivariate time series (Gungor et al., 2025), with practical applications in fraud detection systems like AWS Fraud Detector (web, ).
@@
  • Biological Classification. OOD detection identifies unseen species in biological classification. DNA barcodes enhance image-based classification through re-ranking, improving detection accuracy (Arslan et al., 2024).
@@
  • Spiking Neural Networks (SNN). SNNs’ sparse representations challenge OOD detection. Recent methods combine feedforward learning with distance-based scoring (Terres-Escudero et al., ). The N-MNIST dataset (Orchard et al., 2015) is commonly used for evaluation, with OOD samples generated through diverse event patterns.
  • Astronomical Imaging. OOD detection distinguishes between simulated and real galaxy images. Bayesian comparison and sparse variational autoencoders (Zhou et al., 2024) uncover subtle differences.
@@
  • Edge Computing. Edge computing requires efficient, privacy-preserving OOD detection. SecDOOD (Li et al., 2025) uses cloud-device architecture and HyperNetwork generation to avoid device-side backpropagation. CIFAR-10 and Tiny ImageNet serve as evaluation datasets.
The deployment of OOD detection in high-stakes applications introduces considerations beyond predictive accuracy, spanning regulatory compliance, operational safety, and algorithmic fairness. For compliance, OOD detection supports safety standards like ISO 21448111https://www.iso.org/standard/77490.html by identifying OOD scenarios, informing release decisions and enabling in-operation validation via shadow-mode monitoring (Henriksson et al., 2023). For safety, approaches (Henriksson et al., 2023; Lin et al., 2025) are advancing from runtime input flagging (Henriksson et al., 2023) toward monitoring system-level properties for predictive safety assurance with formal guarantees (Lin et al., 2025). For fairness, mitigating bias against minority subgroups involves developing fairness-aware algorithms (Shekhar et al., 2021) and constructing unbiased benchmarks that control for semantic ambiguity (Inkawhich et al., 2022).
@@ 7.1. Better methodologies of OOD detection
Meta-Learning Adaptation. Addressing the challenge of adapting to new data distributions at test time, meta-learning has gained attention for its ability to facilitate efficient OOD detection in novel scenarios. Recent work also highlights the importance of developing better sampling strategies to more effectively leverage outlier information during training (Ming et al., 2022a).
Model Selection. Choosing an appropriate OOD detection method for a specific task is challenging, as performance varies across data distributions. Automated frameworks like MetaOOD (Qin et al., 2025) use meta-learning for zero-shot selection, while DSDE (Geng et al., 2024) and score-combining methods (Reehorst and Schniter, 2025) apply proportion estimation and likelihood ratio tests to improve ensemble and score fusion. Notably, recent studies show that partially trained models can also be effective for OOD detection (Montazeran and Köthe, 2025), highlighting new possibilities for leveraging model dynamics.
Active Learning. Active learning prioritizes labeling informative data points for efficiency in data-scarce settings. In OOD detection, it focuses on ambiguous or uncertain samples to improve ID and OOD discrimination. SISOM (Schmidt et al., 2024) combines active learning and OOD detection using feature space distances.
Continual and Incremental Learning. Continual and incremental learning enable OOD detection models to adapt to streaming data while preserving previous knowledge. Recent work includes unsupervised OOD detection (Doorenbos et al., 2024), OpenCIL benchmark (Miao et al., 2024b), and hierarchical two-sample tests (Liu et al., 2025d).
@@ 7.2. More practical scenarios of OOD detection
Federated Environments. Federated learning involves training models across decentralized clients, each with its own data distribution (Guo et al., ; Kuang et al., 2024). This distributional variability presents unique OOD detection challenges. Methods like FOOGD (Liao et al., 2024) address these issues by enhancing robustness to both covariate and semantic shifts in federated settings.
Multi-Modal Detection. With the increasing prevalence of multi-modal data, OOD detection has expanded beyond single modalities. MultiOOD (Dong et al., 2024) established a comprehensive benchmark with algorithms like A2D and NP-Mix. Recent works (Gao et al., 2024) have advanced cross-modal alignment techniques.
Multi-Label OOD Detection. Real-world applications often involve multi-label data, where each instance may belong to multiple categories. Detecting OOD samples in these scenarios is challenging due to increased dimensionality and label correlations. Recent approaches (Mei et al., 2024) leverage uncertainty modeling and subjective logic to improve robustness in multi-label OOD detection.
Open-Vocabulary Scenario. Traditional approaches often assume access to all ID categories, an assumption that does not hold in open-vocabulary contexts. Recent work (Li et al., 2024c) explores learning migratable negative prompts, enabling OOD detection even when only a subset of ID labels is available, thus paving the way for more generalizable models.
@@ 7.3. New applications of OOD detection
Human-in-the-Loop Systems. Human expertise integration in OOD detection is crucial for safety-critical applications. Recent frameworks (Vishwakarma et al., 2024; Bai et al., 2024b, c) leverage human feedback to optimize detection thresholds and guide annotation.
Web Image Scraping. OOD detection is also finding novel applications in automating web image scraping. For example, zero-shot ID detection (Miyai et al., 2023c) classifies images as ID if they contain relevant objects, offering a fresh perspective on leveraging OOD detection for large-scale, real-world data acquisition.
@@ 9. Acknowledgements
We thank the editor and reviewers for their constructive feedback, and Zhen Jia and Yongcan Yu for their assistance with the manuscript. This work was supported by the National Natural Science Foundation of China (62276256, U2441251), the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and the Young Scientists Fund of the State Key Laboratory of Multimodal Artificial Intelligence Systems (ES2P100117).